2. The strength of the uniform electric field between two parallel conducting plates separated by 4.00 cm is 7.50×104 V/m . a) What is the potential difference between the plates?  b) The plate with the lowest potential is taken to be at zero volts. What is the potential 1.00 cm from that plate (and 3.00 cm from the other)?  


1
Expert's answer
2022-03-29T12:28:08-0400

(a) the potential difference

"V=Ed\\
=7.5*10^4\:\rm V/m\times0.04\: m=3000\; V"

(b) the potential 1.00 cm from zero potential plate

"V_1=Ed_1\\
=7.5*10^4\:\rm V/m\times0.01\: m=750\; V"